# Conceptualization, Methodology, Writing – original draft, Software, Writing – review & editing

### v2

@@ 5.2 Segment-level
The HBSGRP model for segment-level V–V near misses introduces random parameters within a hierarchical framework (Table 7). Allowing the intercept and three time-varying vehicle-based measures, relative speed, relative deceleration, and relative distance, to vary across segments (See Fig. 5.2). This specification acknowledges that longitudinal dynamics in uninterrupted midblock contexts are shaped not only by vehicle behavior but also by geometric features such as lane configuration, median presence, and driveway density. All remaining covariates, including relative acceleration, jerk, heading and steering differences, traffic volume, lane attributes, and access density, were modeled as fixed effects due to their comparatively stable influence across segments. The HBSGRP specification produces substantial improvements in model fit over the HBSFP baseline, confirming the value of partial pooling in capturing heterogeneity across the nine directional segments analyzed.
For the location parameter ($\mu$), random effects highlight systematic heterogeneity across the nine directional segments. Relative speed exerts a significant positive effect, confirming that higher closing speeds consistently compress available 2D-TTC and elevate near-miss severity. However, the posterior distributions in Fig. 5.2 reveal that the magnitude of this effect varies substantially across segments. Segments with uninterrupted, high-speed flows (e.g., S11) show stronger effects, while others with denser access points (e.g., S6, S7 and S13) exhibit wider credible intervals. Relative deceleration also carries a negative sign, though with more overlap in its credible intervals across segments, suggesting that emergency braking may sometimes mitigate severity in smoother-flow environments, while in higher-speed segments it intensifies near-misses. Notably, in some lower-sample segments, credible intervals for the location intercept approach or cross zero, reflecting weaker estimation stability under sparse BM events. Relative distance displays a consistently negative effect across segments, meaning that shorter headways reliably reduce 2D-TTC, yet its posterior variation underscores differences in spacing behavior along divided versus undivided facilities.
@@
Among the fixed effects, relative acceleration shows a significant negative coefficient in both models, implying that acceleration phases are systematically linked to higher near-miss severity on segments. This contrasts with intersection contexts, where acceleration often reflected flow resumption. Here, aggressive acceleration during overtaking or merging maneuvers reduces spacing and increases risk. Jerk is also negative, reinforcing the role of abrupt longitudinal control in producing critical outcomes. Lateral control indicators—heading and steering differences—show positive signs, suggesting that lateral adjustments modestly increase 2D-TTC where lane-changing flexibility exists. Driveway density exhibits a strong positive effect on $\mu$, which may appear counterintuitive, but is likely explained by anticipatory driver caution in access-rich environments. Conversely, undivided medians sharply reduce $\mu$, indicating that exposure to opposing flows creates consistently severe interactions.
@@
For the scale parameter ($\sigma$), the random effects for relative speed, deceleration, and distance are all negative, showing that extreme dynamics compress outcome variability, leading to uniformly critical near misses. The plots confirm this: high-speed and short-spacing conditions cluster tightly around lower 2D-TTC values, with narrow posteriors in segments such as S12 and S15. In contrast, segments like S10 display wider CI for random scale effects, again reflecting limited BM sample sizes. Fixed effects add further nuance: jerk and heading difference increase scale, suggesting that volatile acceleration or lateral corrections introduce heterogeneity in outcomes, while lane number and lane width reduce scale, indicating that greater maneuvering space fosters more predictable near-misses. Driveway density increases dispersion, consistent with the variability introduced by frequent merging and crossing maneuvers. Median type again exerts a strong negative influence, reinforcing that the absence of raised medians compresses severity outcomes into consistently critical ranges.
@@
Finally, the shape parameter ($\xi$) remains negative and significant across segment groups, confirming a bounded upper tail consistent with the Weibull domain. This indicates physical and behavioral limits to maximum 2D-TTC values during critical segment near-misses, constrained by speed environment, geometry, and driver reaction times. Posterior variation in the shape intercept suggests that undivided segments experience steeper extremal limits, while divided facilities allow marginally greater spread in outcomes. Overall, the segment-level results highlight the primacy of longitudinal dynamics in shaping corridor risk, with acceleration, braking, and headway dominating near-miss severity, while roadway design elements such as medians, lane width, and access density govern both severity and variability. The superior fit of the HBSGRP model underscores the importance of capturing these segment-specific dynamics, as fixed-parameter structures fail to account for the spatially contingent influence of speed and spacing across midblock contexts.
@@
The HBSGRP specification for segment-level V–I near-misses provides a detailed view of how longitudinal dynamics and roadway design influence the severity and variability of extreme 2D-TTC outcomes (Table 8, Fig. 5.2). In this model, the intercept and relative distance are treated as group-level random effects in both the location and scale components of the HBSGRP model, allowing for spatial heterogeneity in how spacing between interacting vehicles shapes COR across segments. All other predictors—including relative speed, acceleration, deceleration, jerk, heading difference, steering angle difference, traffic volume, lane attributes, driveway density, median type, and left-lane presence, are modeled as fixed effects, reflecting their expected stability across segment environments.
@@
For the location parameter, shorter relative distances strongly reduce 2D-TTC values, confirming that closer spacing consistently produces more critical near-miss outcomes. The random effect for distance further indicates that influence varies by segment group, suggesting sensitivity to local geometric or operational conditions. Relative speed displays a notable divergence: while the fixed-parameter benchmark shows a negative coefficient, the grouped model produces a small but positive estimate, implying that in certain contexts, higher speeds are associated with marginally longer 2D-TTC values. This reversal highlights that segment-level heterogeneity in speed environments, such as free-flow arterials versus constrained multilane facilities, can fundamentally alter the role of speed in shaping severity. Relative acceleration emerges as a significant positive effect in the grouped specification, suggesting that acceleration phases along segments may often represent clearance or recovery maneuvers rather than heightened near-miss risk. In contrast, relative deceleration maintains a negative coefficient across both models, reflecting the reality that braking during close-following conditions is tied to critical reductions in 2D-TTC, likely triggered by sudden evasive actions.
@@
Several geometric and operational features are also important for location. Heading and steering differences exhibit significant positive coefficients, indicating that lateral adjustments yield small gains in clearance and mitigate near-miss severity. Volume, lane number, and lane width also emerge as positive contributors in the HBSGRP specification, suggesting that higher-capacity, wider facilities afford drivers greater maneuvering space, leading to slightly larger 2D-TTC values. The effect of driveway density, however, is particularly striking: while insignificant in the fixed-parameter model, the grouped specification produces a large negative estimate with a wide CI. This instability suggests that driveway effects are highly localized—segments with dense access points may concentrate more severe near-miss outcomes, but the data are sparse and sensitive to segment-specific characteristics. Median type is also significant, with undivided facilities consistently producing lower 2D-TTC values, reflecting heightened exposure to opposing traffic flows. The left-lane variable shows a sign reversal, with the fixed model suggesting higher severity in interior lanes and the grouped model indicating greater severity in outer lanes. This again illustrates heterogeneity across segment groups, where the positional context of a lane relative to roadside boundaries and merging activity shifts the locus of risk.
@@
The scale parameter provides complementary insights into the dispersion of near-miss severity. The random effect for relative distance is positive, showing that tighter spacing increases not only severity but also variability in outcomes, as different drivers adopt divergent strategies under minimal headway conditions. Relative speed also has a positive association with scale, indicating that higher speeds amplify heterogeneity in outcomes across V–I interactions, unlike in V–V segment near-misses, where speed reduces dispersion. Acceleration, by contrast, has a negative coefficient in the grouped model, suggesting that acceleration phases yield more predictable outcomes, consistent with smoother-flow recovery dynamics. Relative deceleration also reduces variability, implying that braking in near-crash contexts converges toward uniformly critical outcomes.
@@
Geometric attributes reveal nuanced effects on variability. Lane width exerts a positive effect, suggesting that while wider lanes modestly reduce average severity, they also introduce greater variability in outcomes by enabling diverse driver responses. The undivided median type strongly increases scale, underscoring that undivided facilities not only heighten the average risk but also broaden the range of severity outcomes, likely due to variation in driver avoidance maneuvers. The left-lane effect is positive in scale, indicating that severity outcomes in these lanes are more heterogeneous, perhaps reflecting the diversity of driver strategies in navigating outer versus inner lanes. Driveway density, although statistically unstable, shows a wide and imprecise effect that suggests variability is strongly contingent on site-specific access management patterns.
@@
The shape parameter remains negative and significant across both specifications, confirming that the upper tail of the 2D-TTC distribution is bounded. This supports the Weibull characterization of extremes, implying that even in segment-level V–I interactions, physical and behavioral limits constrain how severe the most critical near-miss events can be. Variation in shape across segment groups highlights the role of local geometry, access density, and operational features in shaping the extremal properties of near-miss distributions.
@@
Together, the segment-level analyses of V–V and V–I near misses underscore that roadway geometry and vehicle dynamics interact to shape both the severity and variability of extreme near-misses. For V–V events, relative speed and distance dominate as key random effects, with speed consistently amplifying severity and distance compressing variability, reflecting deterministic outcomes under close-following conditions. For V–I events, however, the role of speed is more heterogeneous—sometimes protective, sometimes risk-enhancing—depending on the segment context. Access-related variables, particularly driveway density and median type, emerge as critical for V–I outcomes, underscoring the importance of access management and median design in shaping roadside safety. In both near-miss types, the negative and significant shape parameter indicates bounded extremes, reinforcing the physical and behavioral limits to crash risk escalation.
@@ 6 Model validation
The dataset for the study corridor provides precise geolocation for each extreme near-miss event and readily available annual corridor-level observed crashes. The dataset for the study corridor provides precise geolocation for each extreme near-miss event from Eqn. 32. To address this limitation, validation was conducted as an event-level case–control approach, designed to assess whether the model assigns higher predicted probabilities to more severe near-miss events. Posterior exceedance probabilities were computed from Eqn. 30 for severity thresholds ranging from –0.1 to –0.9 seconds, and these probabilities were then used to evaluate classification performance using the area under the receiver operating characteristic curve (ROC–AUC). This metric is particularly well-suited to rare-event prediction because it is independent of a specific threshold.
@@
ROC curves were generated for the four HBSGRP–UGEV variants, covering both V-V, V–I interactions at segment and intersection levels. The results (Figure 6) reveal clear distinctions across near-miss types and roadway contexts. The segment-level V-V model consistently achieved the highest accuracy, with mean AUC values of about 0.89 and very little variation across thresholds. This indicates stable and robust predictive performance, reflecting the ability of grouped random parameters to capture high-speed longitudinal near-miss dynamics along uninterrupted midblock segments. The V-I model at the segment level achieved slightly lower accuracy, averaging around 0.79, and exhibited moderate sensitivity to the choice of severity threshold, a finding consistent with the greater variability in driver responses when vehicles approach roadside boundaries. At intersections, the V-V model demonstrated moderately strong predictive accuracy, with an average AUC of nearly 0.82 and relatively low volatility across thresholds. This indicates that grouped speed and spacing dynamics remain reliable even under more complex signalized and turning conditions. By contrast, the intersection-level V-I model performed weakest, averaging an AUC of approximately 0.75 and showing large fluctuations across thresholds, which reflects the complexity of boundary-related near-misses in intersections where turning trajectories, curb proximity, traffic signals, and pedestrian activity all interact.
Taken together, these findings confirm that the HBSGRP–UGEV framework provides strong and reliable performance for V-V interactions, particularly in segment environments, while predictions for V-I interactions are more sensitive to threshold selection and local variability. This suggests that segment-level V-V estimates are the most operationally reliable for proactive safety monitoring, whereas intersection-level V-I near-misses may require further refinement. Incorporating more detailed information on curb geometry, pedestrian exposure, and high-resolution vehicle–infrastructure proximity measures could improve stability and predictive accuracy in future applications.
@@ 7 Summary and conclusions
This study addresses the persistent challenge of accurately estimating short-term crash occurrence risk (COR) in complex, mixed urban traffic environments, where conventional surrogate safety models often overlook detailed vehicle geometry, roadway context, and site-specific behavioral variability. Existing approaches are limited by oversimplified near-miss representations, a predominant focus on vehicle–vehicle (V–V) interactions, and insufficient integration of spatial heterogeneity in statistical modeling. To overcome these gaps, we propose a high-fidelity, geometry-aware, and statistically robust framework for real-time COR estimation that introduces a novel two-dimensional time-to-collision (2D-TTC) indicator. This metric accounts for vehicle dimensions, dynamic states, and precise geometric positioning, enabling accurate detection of extreme near-miss events from both V–V and vehicle–infrastructure (V–I) interactions.
@@
Extreme events are translated into probabilistic risk estimates through a non-stationary Univariate Generalized Extreme Value (UGEV) framework implemented within a Hierarchical Bayesian Spatial Grouped Random Parameter (HBSGRP) structure. A corridor-based grouping strategy was employed to address event sparsity, pooling near-miss samples across intersections and directional segments while preserving local heterogeneity. Execution of the framework involved extracting high-resolution vehicle trajectories from the Argoverse-2 motion forecasting dataset, integrating them with HD maps and supplementary roadway information, and identifying near-miss events. The model incorporated both vehicle dynamics covariate (relative speed, acceleration, deceleration, jerk, heading difference, steering difference, and traffic volume) and geometric or operational covariates (lane number, lane width, driveway density, median type, turning movement, and vehicle lane position). Grouped random parameters were specified for temporally varying dynamics, relative speed, relative deceleration, and relative distance, to capture their context-dependent effects. For benchmarking, a baseline hierarchical Bayesian fixed-parameter (HBSFP) model was also estimated.
@@
Results demonstrate that HBSGRP consistently outperformed the HBSFP across both intersections and segments, achieving lower DIC, WAIC, and LOOIC values and providing greater estimation stability. Relative speed emerged as the dominant predictor of near-miss severity in intersections, while relative distance was more influential for segment-level V–V events. Deceleration played a critical role in segment-level V–V near-miss, underscoring the importance of longitudinal control. Fixed covariates such as heading difference, steering difference, lane width, driveway density, and median type also showed significant effects, demonstrating the ability of the model to capture both dynamic behaviors and roadway design influences.
@@
Validation results further confirmed model credibility. ROC–AUC analysis showed strong predictive performance, with V–V segment models achieving the highest accuracy (AUC $\approx$ 0.89), followed by intersection V–V ($\approx$ 0.82). V–I models performed reasonably well (segment $\approx$ 0.79; intersection $\approx$ 0.75) but displayed greater sensitivity to severity threshold choice. These findings highlight that while V–V interactions are modeled with high stability, V–I near-misses especially at intersections, remain more complex and variable.
@@
Despite its contributions, the study has several limitations. First, potential bias may arise from the dependence on AV onboard sensors, which may affect traffic state accuracy and near-miss detection. Second, the 2D-TTC formulation, while tractable, simplifies nonlinear vehicle dynamics, which could understate risks in extreme evasive maneuvers. Third, the current framework does not explicitly model near-miss severity levels. Future research should extend the framework to multivariate UGEV models incorporating complementary surrogate measures such as post-collision velocity change ($\Delta V$) and explicitly address vulnerable road user (VRU) near-misses. Broader validation across multiple cities, transferability calibration, and comparison with peaks-over-threshold (POT) GPD methods will further strengthen robustness and generalizability. As autonomous vehicle deployment increases, access to larger and richer trajectory datasets will enhance the statistical power of extreme value analysis and enable more nuanced safety assessments. In practice, the proposed HBSGRP–UGEV framework offers a scalable, data-driven tool for proactive safety monitoring at the corridor level. By enabling early identification of high-risk conditions and supporting real-time traffic management, it provides a pathway toward Vision Zero objectives, reducing fatalities and serious injuries while promoting safe, efficient, and equitable mobility in urban environments.

### v3

@@ 5.1.2 Segment-level
The HBSGRP model for segment-level V–V near misses introduces random parameters within a hierarchical structure (Table 7). Specifically, the intercept and three time-varying vehicle dynamics—relative speed, relative deceleration, and relative distance—were allowed to vary across segments (Fig. 5.1.2). This specification recognizes that longitudinal interactions in uninterrupted midblock contexts are shaped by both driver behavior and roadway characteristics (lane configuration, median type, driveway density). All remaining covariates—relative acceleration, jerk, heading and steering differences, traffic volume, lane attributes, and access density—were treated as fixed effects given their comparatively stable influence across segments. The HBSGRP model yields substantially better fit than the fixed-effect alternative, confirming the value of partial pooling for capturing heterogeneity across the nine directional segments.
@@
For the location parameter ($\mu$), the random effects reveal meaningful differences across segments that align with expected driving dynamics. Relative speed shows a consistently positive effect, confirming that faster closing speeds sharply reduce 2D-TTC and intensify near-miss severity. The magnitude of this effect, however, varies across roadway contexts Fig. 5.1.2. Segments with uninterrupted high-speed flows (e.g., S11) show the strongest impacts, while access-rich segments (e.g., S6, S7, and S13) display wider credible intervals, reflecting greater variability in interactions. Relative deceleration carries a negative sign overall, but its effect is less stable, suggesting that braking sometimes helps dissipate conflicts in smoother-flow environments, whereas in higher-speed segments the same braking often signals critical near-miss events. Relative distance exhibits a consistently negative association, indicating that shorter headways reliably elevate severity by restricting the available buffer for evasive maneuvers. The posterior variation across segments suggests that spacing behavior differs between divided and undivided facilities. Finally, in some segments with sparse block maxima samples, the intercept’s credible intervals approach zero, highlighting weaker estimation stability under limited data conditions.
@@
Among the fixed effects, relative acceleration shows a significant negative association with $\mu$, indicating that acceleration phases are systematically linked to more severe near-misses on segments. This pattern contrasts with intersections, where acceleration often reflects flow resumption under green signals. Along segments, however, aggressive acceleration during overtaking or merging reduces spacing and elevates risk. Jerk also exhibits a negative effect, underscoring how abrupt longitudinal control contributes to critical outcomes. Lateral control indicators—heading and steering differences—display positive coefficients, suggesting that lateral adjustments provide modest increases in 2D-TTC where lane-changing flexibility is available. Driveway density is positively associated with $\mu$, a result that may seem counterintuitive but likely reflects anticipatory caution by drivers in access-rich environments. In contrast, undivided medians show a strong negative effect, confirming that exposure to opposing traffic flows produces consistently more severe interactions.
@@
For the scale parameter ($\sigma$), the random effects for relative speed, deceleration, and distance are all negative, indicating that extreme longitudinal dynamics compress outcome variability and yield uniformly critical near-miss conditions. This pattern is evident in Fig. 5.1.2, where high-speed and short-spacing conditions cluster tightly around lower 2D-TTC values, producing narrow posteriors in segments such as S12 and S15. By contrast, segments like S10 display wider credible intervals, reflecting weaker estimation stability under sparse BM samples. Among the fixed effects, jerk and heading difference increase $\sigma$, suggesting that volatile acceleration or lateral corrections introduce greater heterogeneity in near-miss outcomes. In contrast, lane number and lane width reduce $\sigma$, implying that additional maneuvering space fosters more predictable interactions. Driveway density increases dispersion, consistent with the variability created by frequent merging and crossing maneuvers. Median type again shows a strong negative effect, reinforcing that undivided facilities compress severity outcomes into consistently critical ranges.
@@
Finally, the shape parameter ($\xi$) remains negative and statistically significant across segment groups, confirming a bounded upper tail consistent with the Weibull domain. This finding indicates physical and behavioral constraints on extreme near-miss severity, governed by speed environment, roadway geometry, and driver reaction capacity. Overall, the segment-level results emphasize the central role of longitudinal dynamics—acceleration, braking, and headway—in determining near-miss severity, while roadway features such as medians, lane width, and access density primarily shape variability. The strong performance of the HBSGRP model underscores its value in capturing these segment-specific dynamics, which fixed-parameter structures fail to represent, particularly in accounting for the spatially contingent effects of speed and spacing in midblock contexts.
@@
The HBSGRP model for segment-level V–I near-misses offers insights into how longitudinal dynamics and roadway design jointly shape the severity and variability of extreme 2D-TTC outcomes (Table 8, Fig. 5.1.2). In this model, the intercept and relative distance are specified as group-level random effects in both the location and scale components, capturing spatial heterogeneity in how spacing influences near-miss severity across segments. All other covariates, including relative speed, acceleration, deceleration, jerk, heading difference, steering angle difference, traffic volume, lane attributes, driveway density, median type, and left-lane presence—are modeled as fixed effects to represent their more consistent contributions across segment environments.
@@
For the location parameter, reduced clearance to roadway boundaries exerts a strong negative effect on 2D-TTC, confirming that closer spacing to infrastructure consistently produces more critical near-miss outcomes. The random effect for distance further indicates that this influence varies across segment groups, reflecting sensitivity to local geometric and operational conditions such as median presence or driveway density. Relative speed shows a notable divergence: while the fixed-parameter model suggests higher speeds intensify severity, the grouped specification produces a small positive estimate, implying that in some contexts faster travel may coincide with greater clearance from boundaries (e.g., drivers maintaining lane center on free-flow arterials). This reversal highlights how heterogeneous speed environments fundamentally alter the role of speed in shaping V–I severity. Relative acceleration is positive and significant, suggesting that acceleration phases often represent clearance from roadside constraints (e.g., moving away from curbs or barriers) rather than heightened risk. By contrast, relative deceleration retains a negative effect across both models, indicating that braking near boundaries is typically associated with abrupt evasive actions, sharply reducing 2D-TTC and producing more critical near-miss conditions.
@@
Several geometric and operational features also shape V–I near-miss severity at the segment level. Heading and steering differences exhibit significant positive coefficients, indicating that lateral adjustments toward or away from boundaries provide modest clearance gains and mitigate severity. Traffic volume, lane number, and lane width also emerge as positive contributors in the HBSGRP specification, suggesting that wider, higher-capacity segments allow greater separation from roadside features, producing slightly larger 2D-TTC values. The effect of driveway density, however, is particularly notable: while insignificant in the fixed-parameter model, the grouped specification yields a large negative estimate with a wide CI. This suggests that dense roadside access points intensify boundary conflicts by forcing frequent merging and turning maneuvers, though the data remain sensitive to localized characteristics. Median type is also significant, with undivided facilities consistently producing lower 2D-TTC values, reflecting heightened exposure to opposing flows and limited recovery space near boundaries. The left-lane variable shows a sign reversal, with the fixed model suggesting higher severity in interior lanes and the grouped model indicating greater severity in outer lanes. This pattern highlights heterogeneity across segment groups, where lane position relative to curbs, medians, or roadside access alters the likelihood of boundary-related near misses.
@@
The scale parameter provides complementary insights into the variability of boundary-related near misses. The random effect for relative distance is positive, indicating that reduced clearance to roadside features not only increases severity but also broadens variability, as drivers adopt divergent strategies when negotiating minimal spacing with curbs, medians, or barriers. Relative speed also shows a positive association with scale, suggesting that higher speeds amplify heterogeneity in boundary interactions—contrasting with V–V near misses, where speed reduces dispersion—because drivers respond differently to infrastructure proximity under fast-moving conditions. Acceleration, by contrast, has a negative coefficient in the grouped model, implying that acceleration phases near boundaries often reflect smoother clearance maneuvers, yielding more predictable outcomes. Relative deceleration also reduces variability, indicating that braking close to infrastructure compresses outcomes into uniformly critical events.
@@
Geometric attributes further shape variability. Lane width exerts a positive effect, showing that while wider lanes provide greater average clearance, they also increase variability by enabling a broader range of lateral positioning relative to boundaries. The undivided median type strongly increases scale, underscoring that the absence of a raised median not only heightens risk but also expands outcome variability due to heterogeneous driver avoidance maneuvers. The left-lane variable is also positive, indicating that boundary proximity in outer lanes produces more heterogeneous near-miss outcomes compared to interior lanes. Finally, driveway density, though statistically unstable, suggests that variability is strongly site-dependent, as frequent access points create unpredictable driver–infrastructure interactions.
@@
The shape parameter remains negative and significant across both specifications, confirming that the upper tail of the 2D-TTC distribution is bounded. This supports the Weibull domain for extremes and indicates that even in V–I segment interactions, physical and behavioral limits constrain how severe the most critical boundary-related near-misses can become. Variation in the shape parameter across segment groups underscores the role of local geometry, access density, and operational conditions in shaping the extremal behavior of these events.
@@
Segment-level analyses show that roadway geometry and vehicle dynamics jointly shape near-miss severity and variability. For V–V events, relative speed and distance dominate—speed amplifies severity while distance compresses variability, producing more deterministic outcomes under close-following conditions. For V–I events, the effect of speed is heterogeneous, sometimes reducing severity when clearance to boundaries is maintained, but elevating it when high speeds bring vehicles closer to curbs or medians. Access-related factors, particularly driveway density and median type, are critical for V–I outcomes, underscoring the role of access management and median design in roadside safety. Across both interaction types, the negative shape parameter indicates bounded extremes, reflecting physical and behavioral limits on crash risk escalation.
@@ 6 Model validation
The study corridor dataset provided precise geolocation for each extreme near-miss event (Eqn. 32) along with annual corridor-wide crash totals. Although annual crash data offer a useful aggregate benchmark, they lack precise geolocation of individual crashes and therefore cannot be directly matched to near-miss predictions at the event level. To address this limitation, validation was instead designed as an event-level case–control exercise, assessing whether the model assigns higher exceedance probabilities to more severe near-miss events. Posterior exceedance probabilities were computed (Eqn. 30) across severity thresholds from –0.1 to –0.9 seconds, and classification performance was evaluated using ROC–AUC, a threshold-independent metric well suited to rare-event prediction.
@@
ROC curves were generated for the four HBSGRP–UGEV variants, covering both V-V and V-I interactions at the segment and intersection levels. The results (Figure 6) reveal clear distinctions across near-miss types and roadway contexts. The V–V segment model achieved the highest predictive accuracy (AUC $\approx$ 0.89), reflecting the influence of dynamic variability in driver behavior. Relative speed, deceleration, and distance emerged as significant random parameters, capturing heterogeneity in conflict severity across segments. At the same time, fixed effects such as heading difference, steering difference, jerk, lane width, and driveway density contributed additional explanatory power, indicating that both dynamic interactions and roadway design shape V–V near-miss risks. The V–I model at the segment level achieved moderately strong accuracy (AUC $\approx$ 0.79) but displayed greater sensitivity to severity thresholds. Only relative distance emerged as a significant random parameter, while other vehicle dynamics (e.g., speed, deceleration, etc.) and roadway features (e.g., lane width, driveway density, median type, and etc.) contributed as fixed effects. This suggests that V–I risks are shaped primarily by localized conditions, with dynamic variability playing a comparatively smaller role than in V–V interactions.
At intersections, the V–V model achieved moderately strong accuracy (AUC $\approx$ 0.82) with low threshold sensitivity. Random parameters for relative speed and relative distance captured cross-location heterogeneity in approach and gap-acceptance conditions, while fixed effects— including deceleration/acceleration, steering and heading differences, jerk, and a left-turn indicator—explained lateral maneuvering and control effects. Together, these terms allow the model to represent both longitudinal and lateral conflict dynamics under signalized conditions, yielding stable performance. In contrast, the intersection-level V-I model performed weakest (AUC $\approx$ 0.75) and showing large fluctuations across thresholds, which reflects the complexity of boundary-related near-misses in intersections where turning trajectories, curb proximity, traffic signals, and pedestrian activity all interact.
Taken together, the results demonstrate that the HBSGRP–UGEV framework yields robust and consistent predictive accuracy for V–V interactions, particularly along corridor segments, where conflict dynamics are more regular and less influenced by intersectional variability. By contrast, V–I models—especially at intersections—showed lower stability and greater sensitivity to threshold specification, reflecting the complexity of boundary-related near-miss shaped by turning flows, curb geometry, and multimodal activity. These findings suggest that segment-level V–V estimates are the most operationally reliable for proactive monitoring, while intersection-level V–I models may require further refinement. Incorporating richer representations of curb design, pedestrian exposure, and high-resolution V-I proximity measures could strengthen stability and predictive performance in future applications.
@@ 7 Summary and conclusions
This study addresses the persistent challenge of estimating short-term crash occurrence risk (COR) in urban corridor-wide traffic environments, where conventional surrogate safety models often overlook detailed vehicle geometry, roadway context, and site-specific behavioral variability. Existing approaches have typically focused on vehicle–vehicle (V–V) interactions and neglected vehicle–infrastructure (V–I) conflicts, limiting their applicability at the corridor scale. To overcome these gaps, we introduced a geometry-aware two-dimensional time-to-collision (2D-TTC) indicator that incorporates vehicle dimensions, dynamic states, and roadway boundaries, enabling the detection of extreme near-miss events from both V–V and V–I interactions.
@@
Extreme events were modeled through a nonstationary Univariate Generalized Extreme Value (UGEV) distribution embedded within a Hierarchical Bayesian Spatial Grouped Random Parameter (HBSGRP) framework. This structure allowed partial pooling across intersections and directional segments while preserving local heterogeneity, addressing data sparsity and contextual variation. High-resolution vehicle trajectories from the Argoverse-2 dataset were integrated with HD maps to extract near-miss events. Vehicle dynamics (relative speed, acceleration, deceleration, jerk, heading and steering differences, and traffic volume) and roadway features (lane width, driveway density, median type, turning movements, and lane position) were incorporated as covariates, with grouped random parameters specified for key dynamic variables. Grouped random parameters were specified for temporally varying dynamics, relative speed, relative deceleration, and relative distance, to capture their context-dependent effects. For benchmarking, a baseline hierarchical Bayesian fixed-parameter (HBSFP) model was also estimated.
@@
Results showed that the HBSGRP model consistently outperformed its fixed-parameter counterpart, yielding lower DIC, WAIC, and LOOIC values and demonstrating greater stability. Random-parameter vehicle dynamics revealed that relative speed, deceleration, and distance dominated V–V near misses at the segment level, whereas only relative distance was significant for V–I near misses. At intersections, V–V near misses were dominated by relative speed and distance, while no dynamic variables were significant for V–I near misses. Fixed covariates such as heading difference, steering difference, lane width, driveway density, and median type also showed significant effects, demonstrating the ability of the model to capture both dynamic behaviors and roadway design influences.
@@
Validation results further confirmed model credibility. ROC–AUC validation confirmed robust predictive accuracy, with V–V segment models performing best (AUC $\approx$ 0.89), followed by intersection V–V ($\approx$ 0.82)), while V–I models showed moderate performance ($\approx$ 0.79 for segments, $\approx$ 0.75 for intersections). These findings indicate that while V–V risks can be reliably modeled, V–I interactions, particularly at intersections, remain more variable and complex.
@@
While the proposed framework advances corridor-wide COR modeling, several aspects warrant further development. First, reliance on AV onboard sensors may introduce measurement bias in traffic state estimation and near-miss detection. Second, the 2D-TTC formulation simplifies nonlinear vehicle dynamics, which could understate risks during extreme evasive maneuvers. Third, the framework does not yet capture near-miss severity stratification. Future research offers opportunities to enrich this framework. Extending it to multivariate HBSGRP-UGEV models with complementary surrogate measures such as post-collision velocity change ($\Delta V$) would provide deeper insights into crash severity dynamics. Incorporating vulnerable road user (VRU) near-misses, validating across diverse urban networks, calibrating for transferability, and benchmarking against GPD methods will further enhance robustness and generalizability. As AV deployment expands, increasingly rich trajectory datasets will strengthen the statistical power of HBSGRP–UGEV and open new opportunities for nuanced, multimodal safety assessments. Within this evolving data landscape, the proposed framework provides a scalable and adaptable foundation for proactive corridor-wide safety monitoring. By enabling early identification of high-risk conditions and supporting real-time traffic management, the framework aligns with the broader Vision Zero agenda, advancing the long-term goal of safer, more efficient, and equitable urban mobility.